OpenStudy (bbb911):

____2a____ + _____4a_____ a + 1 a2 - 1 Add and simplify will give medal and will fan you :)

OpenStudy (anonymous):

is it \[\frac{ 2a }{ a+1 } + \frac{ 4a }{ a^2-1 }\]

OpenStudy (anonymous):

?

OpenStudy (bbb911):

yes

OpenStudy (anonymous):

get the common factor...

OpenStudy (bbb911):

which is 2a and 4a right?

OpenStudy (anonymous):

no... the LCD...

OpenStudy (bbb911):

a + 1

OpenStudy (anonymous):

but u have a^2-1 right.. first factor that and u will get (a+1)(a-1).. so ur LCD is (a+1)(a-1)... u get me..?

OpenStudy (the_fizicx99):

\(\ \sf \dfrac{2a}{a + 1} + \dfrac{4a}{a^2 - 1} \) a^2 - 1 ==> (a - 1)(a + 1) (2a)/(a + 1) + (4a)/(a - 1)(a + 1) Get a common denominator, multiply both sides by (a - 1)(a + 1) \(\ \sf \dfrac{(2a) \times (a - 1)}{(a + 1)\times(a - 1)} + \dfrac{4a}{(a + 1)(a - 1)} \) Now add them, \(\ \dfrac{2a(a - 1)}{(a + 1)(a - 1)} + \dfrac{4a}{(a + 1)(a - 1)} \implies \dfrac{2a(a - 1) + 4a}{(a + 1)(a-1)} \implies \dfrac{a( 2(a - 1) + 4)}{(a - 1)(a + 1)}\) Simplify 2(a - 1), 2a - 2 \(\ \sf \dfrac{a(2a - 2 + 4)}{(a - 1)(a + 1)} \implies \dfrac{a(2a + 2)}{(a - 1)(a +1)} \) This can be further factored. 2a + 2 ==> 2(a + 1) \(\ \sf \dfrac{a(2(a + 1))}{(a + 1)(a - 1)} \) This cancels out, \(\ \sf \dfrac{a(2\cancel{(a + 1)}}{\cancel{(a + 1)}(a - 1)} \) \(\ \dfrac{a(2)}{(a - 1)} \implies \) \(\ \sf \dfrac{2a}{(a - 1)} \)

OpenStudy (bbb911):

I can see it clearly now :) both of you

OpenStudy (the_fizicx99):

~_^